# Leveraging Machine Learning for High-Dimensional Option Pricing within the Uncertain Volatility Model

### v1

@@ 5 Numerical Experiments
In this Section we report some numerical results in order to investigate the effectiveness of the proposed GTU algorithm for pricing European options in high dimension.
@@
Parameters employed for the tests are shown in Table 1. The options chosen for these tests, along with their numerical parameters, are identical to those analyzed by Guyon and Labordère [6]. This allows us to directly compare our results with theirs.
The algorithm has been implemented in MATLAB and computations have been preformed on a server which employs a 2.40 GHz Intel Xenon processor (Gold 6148, Skylake) and 20 GB of RAM. To better evaluate the computational time of the algorithm, parallelization was not used for solving the optimization problems (4.3).
@@
$T$ Maturity $1.0$ $M$ Number of branches for GPR-Tree variable
@@ 5.1.1 Outperformer option
The Outperformer option is a financial derivative that pays a payoff equal to the positive part of the difference between two asset values at maturity. Specifically, the payoff is given by
$\Psi\left(\mathbf{S}_{T}\right)=\left(S_{T}^{2}-S_{T}^{1}\right)_{+}.$ (5.1)
Thus, the valuation of this option is a 2-dimensional problem, which can be further simplified to a problem in dimension 1 when $\rho_{1,2}\leq 0$ by employing the change of numèraire technique as described by Guyon and Labordère [6]. Thus, although the dimension of this problem is not large, it is interesting in that we have a reliable benchmark (computed for example with the PDE method by Windcliff et al. [13]) as well as the results reported in the paper by Guyon and Labordère [6].
Table 2 provides the numerical results. The results obtained with GTU closely match the benchmark, even for small values of $N$ and $P$. The computational time is minimal.
$N=16$ $N=32$ $N=64$
$\rho_{1,2}$ $P$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $GL$ $BM$
@@
Table 2: Outperformer option. The column labeled “GL” presents the results from [6] for the same option, while the column labeled “BM” shows the benchmark computed by using the change of numéraire technique. Values in brackets indicate the computational times, measured in seconds.
@@ 5.1.2 Outperformer spread option
Similar to the Outperformer option discussed in Subsection 5.1.1, the Outperformer spread option is another example of 2-dimensional option which has been investigated by Guyon and Labordère [6] and for which, when $\rho_{1,2}\leq 0$, a benchmark is available by using the change of numèraire technique. Specifically, the payoff of such an option reads out
@@
Results are presented in Table (3) are very close to the benchmark.
$N=32$ $N=64$ $N=128$
@@
Table 3: Outperformer spread option. The column labeled “GL” presents the results from [6] for the same option, while the column labeled “BM” shows the benchmark computed by using the change of numéraire technique. Values in brackets indicate the computational times, measured in seconds.
@@ 5.1.3 Geo-Call spread
This case is a generalization of the UVM model in that the dividend is not constant but changes in relation to the value of volatility. The binomial tree method proposed by Avellaneda et al. [1] for the 1d UVM model can be used to calculate the price in this model, taking care to change the dividend in relation to volatility. In this regard, it should be noted that while in the one-dimensional UVM model the optimal $\sigma$ value is always $\sigma^{\min}$ or $\sigma^{\max}$ (this feature is known as “the bang bang condition”), in this model, characterized by a variable dividend yield $\hat{\eta}$, this property no longer applies. Therefore, at each node of the tree, the choice of optimal correlation is to be pursued not merely by comparing the objective function at the extreme values of $\sigma$, but by optimising the option price via an appropriate algorithm, such as the Sequential Quadratic Programming algorithm.
The results are shown in Table 4. As shown, the values obtained for varying $N$ and $P$ are very close to the benchmark for the considered values of $d$. The largest deviation is observed for $d=5$, with a difference of $0.04$ between the GTU price and the benchmark. However, this difference is minor, being less than $0.5\%$ of the benchmark. Additionally, we observe that the computational time increases linearly with the number of time steps $N$ and approximately linearly with the number of points $P$. This is particularly interesting since the cost for the GPR regression is cubic in $P$, indicating that most of the computational time is spent on solving optimization problems rather than on the regression itself. Lastly, the computational time grows exponentially with problem size, as the number of values to be evaluated with a tree step is $2^{d}$. Despite this, we are able to handle high dimensions, such as $d=10$, with reasonable computational effort.
$N=16$ $N=32$ $N=64$ $BM$
@@
Table 4: Geo-Call spread, with $\rho_{i,j}=0$ and with $M=2^{d}$. Values in brackets represent the computational times, measured in seconds.
To prevent the computational time from increasing exponentially, we can reduce the number of terms in the sum by randomly selecting a sample of $M$ elements. This procedure is justifiable in that, as $M$ approaches $2^{d}$ , the calculated price converges to the value we would get without this simplification. The results, shown in Table 5, indicate that even with just a few thousand samples, the outcomes are very stable. In addition, we can achieve high size of $d=40$ with reasonable computation time.
$N=16$ $N=32$ $N=64$

### v2

@@ 5 Numerical Experiments
In this Section we report some numerical results in order to investigate the effectiveness of the proposed algorithms for pricing European options in high dimension.
@@
The parameters employed in our numerical experiments are summarized in Table 1. Notably, the options considered – including numerical settings and the bounds for the UVM – coincide with those examined by [Guyon and Henry-Labordère, 2011], thereby enabling a direct comparison of the results. In addition to these benchmark configurations, we also investigate some modified scenarios, characterized by wider volatility bounds and increased problem dimensionality, with the aim of evaluating the robustness and scalability of the numerical methods under more challenging conditions. These test cases, both benchmark and modified, provide the basis for a comparative assessment of the two algorithms under varying conditions. The GTU algorithm has been implemented in Matlab, while the NNU algorithm has been implemented in Python. Computations have been preformed on a server which employs a 2.40 GHz Intel Xenon processor (Gold 6148, Skylake) and 20 GB of RAM. To better evaluate the computational time of the algorithms, parallelization was not used in either algorithm. All calculations were performed on CPU (no GPU use).
Remark 6.
We employ the Matlab implementation of the SQP algorithm, using its standard settings. Specifically, the maximum number of iterations is set to $200$, which provides sufficient iterations for convergence in most practical scenarios. The stopping criterion is determined by the tolerance on the optimality conditions, with a default value of $1\cdot 10^{-6}$. This choice ensures a balance between achieving a solution with adequate accuracy for most applications and maintaining computational efficiency.
Remark 7.
Unless stated otherwise, all results presented in this section were obtained using a feedforward NN with two hidden layers, each consisting of 32 neurons. The activation function is ReLU, denoted by $\phi=\mathrm{ReLU}$, and the training algorithm used is Adam with default parameters. The number of Monte Carlo simulations for NNU is $M=10^{5}$. Moreover, we set the penalization parameter $\lambda$ to $1$. We tried other combinations of parameters but the results are essentially the same as those presented below.
Remark 8.
For technical reasons, the two algorithms have been implemented on different software platforms: GTU in Matlab and NNU in Python. As a result, we do not attempt a detailed comparison of computational times, since such metrics may be influenced by differences in language efficiency, libraries, and hardware usage. Instead, we restrict ourselves to general observations aimed at highlighting only significant or macroscopic differences in computational performance.
@@
$T$ Maturity $1.0$ $M$ Number of branches for GTU variable
$E$ Number of training epochs for NNU variable
@@ 5.1.1 Outperformer option
The Outperformer option is a financial derivative that provides a payoff equal to the positive part of the difference between the values of two assets at maturity. Notably, this type of option is also referred to as an Exchange option and was first introduced by [Margrabe, 1978]. Specifically, the payoff can be expressed as:
$\Psi\left(\mathbf{S}_{T}\right)=\left(S_{T}^{2}-S_{T}^{1}\right)_{+},$ (5.1)
where $\left(x\right)_{+}$ denotes the positive part of $x$, defined as $\max(x,0)$. The valuation of this option constitutes a 2-dimensional problem, which can be reduced to a one-dimensional problem under the condition $\rho_{1,2}\leq 0$. This simplification is achieved by applying the change of numéraire technique, as detailed in [Guyon and Henry-Labordère, 2011]. Thus, although the dimension of this problem is not large, it is interesting in that we have a reliable benchmark (computed for example with the PDE method by [Windcliff et al., 2006]) as well as the results reported in the paper by [Guyon and Henry-Labordère, 2011].
Table 2 reports the numerical results obtained with both the GTU and NNU algorithms. Note that, as for NNU, we also include the $95\%$ margin errors, which remain essentially constant as $E$ varies. In both cases, the outcomes are in close agreement with the benchmark, even for small values of $N$ and $P$ in the GTU method, and for moderate values of $E$ and $P$ in the NNU approach. The computational times for the two algorithms are comparable and remain consistently low. In this low-dimensional setting ($d=2$), GTU appears to outperform NNU, delivering more accurate results for the same computational effort.
To assess the robustness of both algorithms under more challenging market conditions, we also tested them using wider volatility bounds, specifically $\left[\sigma^{\min}_{i},\sigma^{\max}_{i}\right]=[0.1,0.4]$. This configuration amplifies the uncertainty in the underlying dynamics and broadens the admissible set of control variables. The corresponding results are reported in Table 3 and can be directly compared with those in Table 2, which refer to the narrower interval $[0.1,0.2]$. By comparison, we observe that both GTU and NNU continue to perform accurately under increased volatility uncertainty, consistently matching the benchmark across all correlation levels. However, the numerical values in Table 3 exhibit a larger spread, as expected, and the differences between parameter configurations become more pronounced. The NNU method still achieves high accuracy, albeit with a slightly increased margin of error due to the broader range of diffusion inputs. In terms of computational times, both methods remain efficient, with no significant deterioration despite the increased complexity of the problem.
GTU $N=16$ $N=32$ $N=64$
$\rho$ $P$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ GL BM
@@
NNU $N=16$ $N=32$ $N=64$
$\rho$ $E$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ ME GL BM
$-0.5$ $\underset{\left(40\right)}{13.21}$ $\underset{\left(67\right)}{13.62}$ $\underset{\left(121\right)}{13.65}$ $\underset{\left(229\right)}{13.65}$ $\underset{\left(83\right)}{13.14}$ $\underset{\left(144\right)}{13.66}$ $\underset{\left(266\right)}{13.69}$ $\underset{\left(510\right)}{13.70}$ $\underset{\left(156\right)}{13.41}$ $\underset{\left(274\right)}{13.75}$ $\underset{\left(506\right)}{13.78}$ $\underset{\left(968\right)}{13.79}$$\pm 0.13$ $13.77$ $13.75$
00 $\underset{\left(39\right)}{10.81}$ $\underset{\left(66\right)}{11.16}$ $\underset{\left(120\right)}{11.18}$ $\underset{\left(227\right)}{11.19}$ $\underset{\left(88\right)}{11.01}$ $\underset{\left(154\right)}{11.31}$ $\underset{\left(287\right)}{11.34}$ $\underset{\left(553\right)}{11.35}$ $\underset{\left(156\right)}{11.02}$ $\underset{\left(274\right)}{11.21}$ $\underset{\left(506\right)}{11.23}$ $\underset{\left(967\right)}{11.24}$$\pm 0.10$ $11.26$ $11.25$
$0.5$ $\underset{\left(34\right)}{7.64}$ $\underset{\left(58\right)}{7.99}$ $\underset{\left(106\right)}{8.01}$ $\underset{\left(203\right)}{8.02}$ $\underset{\left(75\right)}{7.81}$ $\underset{\left(129\right)}{7.92}$ $\underset{\left(237\right)}{7.94}$ $\underset{\left(452\right)}{7.94}$ $\underset{\left(148\right)}{7.82}$ $\underset{\left(263\right)}{7.99}$ $\underset{\left(495\right)}{8.00}$ $\underset{\left(960\right)}{8.01}$$\pm 0.07$
Table 2: Outperformer option. The column labeled “GL” reports the results from [Guyon and Henry-Labordère, 2011] for the same option, while the column labeled “BM” provides the benchmark computed using the change of numéraire technique. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
GTU $N=16$ $N=32$ $N=64$
$\rho$ $P$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ $125$ $250$ $500$ $1000$ BM
$-0.5$ $\underset{\left(33\right)}{27.19}$ $\underset{\left(54\right)}{27.20}$ $\underset{\left(108\right)}{27.18}$ $\underset{\left(256\right)}{27.17}$ $\underset{\left(62\right)}{27.19}$ $\underset{\left(107\right)}{27.18}$ $\underset{\left(213\right)}{27.16}$ $\underset{\left(527\right)}{27.14}$ $\underset{\left(128\right)}{27.20}$ $\underset{\left(211\right)}{27.19}$ $\underset{\left(426\right)}{27.16}$ $\underset{\left(1120\right)}{27.14}$ $27.10$
00 $\underset{\left(31\right)}{22.35}$ $\underset{\left(51\right)}{22.33}$ $\underset{\left(94\right)}{22.31}$ $\underset{\left(230\right)}{22.31}$ $\underset{\left(57\right)}{22.38}$ $\underset{\left(99\right)}{22.34}$ $\underset{\left(195\right)}{22.32}$ $\underset{\left(494\right)}{22.31}$ $\underset{\left(121\right)}{22.42}$ $\underset{\left(202\right)}{22.38}$ $\underset{\left(427\right)}{22.34}$ $\underset{\left(1078\right)}{22.32}$ $22.27$
$0.5$ $\underset{\left(33\right)}{16.00}$ $\underset{\left(55\right)}{15.94}$ $\underset{\left(97\right)}{15.90}$ $\underset{\left(252\right)}{15.90}$ $\underset{\left(56\right)}{16.06}$ $\underset{\left(104\right)}{15.97}$ $\underset{\left(195\right)}{15.92}$ $\underset{\left(473\right)}{15.90}$ $\underset{\left(105\right)}{16.12}$ $\underset{\left(198\right)}{16.02}$ $\underset{\left(390\right)}{15.94}$ $\underset{\left(985\right)}{15.92}$
NNU $N=16$ $N=32$ $N=64$
$\rho$ $E$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ ME BM
$-0.5$ $\underset{\left(46\right)}{26.46}$ $\underset{\left(75\right)}{27.12}$ $\underset{\left(130\right)}{27.21}$ $\underset{\left(241\right)}{27.23}$ $\underset{\left(80\right)}{25.89}$ $\underset{\left(146\right)}{26.90}$ $\underset{\left(277\right)}{26.99}$ $\underset{\left(542\right)}{27.01}$ $\underset{\left(149\right)}{26.62}$ $\underset{\left(270\right)}{26.98}$ $\underset{\left(512\right)}{27.03}$ $\underset{\left(996\right)}{27.05}$$\pm 0.26$ $27.10$
00 $\underset{\left(32\right)}{21.68}$ $\underset{\left(56\right)}{22.19}$ $\underset{\left(105\right)}{22.26}$ $\underset{\left(202\right)}{22.27}$ $\underset{\left(82\right)}{21.90}$ $\underset{\left(151\right)}{22.23}$ $\underset{\left(289\right)}{22.29}$ $\underset{\left(564\right)}{22.30}$ $\underset{\left(134\right)}{21.89}$ $\underset{\left(242\right)}{22.38}$ $\underset{\left(457\right)}{22.45}$ $\underset{\left(887\right)}{22.47}$$\pm 0.22$ $22.27$
$0.5$ $\underset{\left(37\right)}{15.64}$ $\underset{\left(66\right)}{15.92}$ $\underset{\left(123\right)}{15.96}$ $\underset{\left(238\right)}{15.97}$ $\underset{\left(73\right)}{15.46}$ $\underset{\left(134\right)}{15.74}$ $\underset{\left(256\right)}{15.78}$ $\underset{\left(499\right)}{15.79}$ $\underset{\left(124\right)}{15.44}$ $\underset{\left(222\right)}{15.73}$ $\underset{\left(418\right)}{15.78}$ $\underset{\left(810\right)}{15.79}$$\pm 0.16$
Table 3: Outperformer option with volatility bounds $\left[\sigma_{i}^{\min},\sigma_{i}^{\max}\right]=\left[0.1,0.4\right]$. The column labeled “BM” shows the benchmark computed by using the change of numéraire technique. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
@@ 5.1.2 Outperformer spread option
Similar to the Outperformer option discussed in Subsection 5.1.1, the Outperformer spread option is another example of 2-dimensional option which has been investigated by [Guyon and Henry-Labordère, 2011] and for which, when $\rho_{1,2}\leq 0$, a benchmark is available by using the change of numèraire technique. Specifically, the payoff of such an option reads out
@@
The results, reported in Tables 4 and 5 – for $\left[\sigma^{\min}_{i},\sigma^{\max}_{i}\right]=[0.1,0.2]$ and $\left[\sigma^{\min}_{i},\sigma^{\max}_{i}\right]=[0.1,0.4]$ respectively – confirm that both GTU and NNU remain accurate and reliable even in this more demanding payoff structure.
Overall, both methods succeed in closely matching the benchmark. However, we observe that GTU maintains slightly more stable values across all discretisation parameters and correlation regimes, converging rapidly to the benchmark even with a relatively small number of paths. The NNU method, while slightly more sensitive to the choice of training epochs and exhibiting a marginally wider range of variation in the low-precision regime, still converges consistently as the training size increases. The gap between the two methods slightly widens under increased volatility uncertainty, but remains within acceptable bounds, with NNU still delivering competitive results with limited error margins and moderate computational cost.
GTU $N=32$ $N=64$ $N=128$
@@
NNU $N=32$ $N=64$ $N=128$
$\rho$ $E$ $100$ $200$ $400$ $800$ $1600$ $100$ $200$ $400$ $800$ $1600$ $100$ $200$ $400$ $800$ $1600$ ME GL BM
$-0.5$ $\underset{\left(83\right)}{10.78}$ $\underset{\left(146\right)}{11.17}$ $\underset{\left(272\right)}{11.23}$ $\underset{\left(526\right)}{11.24}$ $\underset{\left(1014\right)}{11.26}$ $\underset{\left(161\right)}{10.63}$ $\underset{\left(281\right)}{11.23}$ $\underset{\left(520\right)}{11.31}$ $\underset{\left(995\right)}{11.33}$ $\underset{\left(1899\right)}{11.34}$ $\underset{\left(296\right)}{10.60}$ $\underset{\left(511\right)}{11.22}$ $\underset{\left(940\right)}{11.33}$ $\underset{\left(1799\right)}{11.35}$ $\underset{\left(3794\right)}{11.37}$$\pm 0.05$ $11.37$ $11.41$
00 $\underset{\left(87\right)}{10.72}$ $\underset{\left(150\right)}{11.18}$ $\underset{\left(279\right)}{11.26}$ $\underset{\left(544\right)}{11.28}$ $\underset{\left(1242\right)}{11.32}$ $\underset{\left(176\right)}{10.61}$ $\underset{\left(309\right)}{11.18}$ $\underset{\left(576\right)}{11.27}$ $\underset{\left(1111\right)}{11.29}$ $\underset{\left(1905\right)}{11.34}$ $\underset{\left(322\right)}{10.56}$ $\underset{\left(573\right)}{11.21}$ $\underset{\left(1071\right)}{11.29}$ $\underset{\left(2068\right)}{11.32}$ $\underset{\left(3813\right)}{11.34}$$\pm 0.05$ $11.39$
$0.5$ $\underset{\left(92\right)}{10.60}$ $\underset{\left(161\right)}{11.03}$ $\underset{\left(299\right)}{11.11}$ $\underset{\left(573\right)}{11.13}$ $\underset{\left(1017\right)}{11.14}$ $\underset{\left(148\right)}{10.64}$ $\underset{\left(256\right)}{11.09}$ $\underset{\left(579\right)}{11.14}$ $\underset{\left(1129\right)}{11.16}$ $\underset{\left(2201\right)}{11.16}$ $\underset{\left(355\right)}{10.47}$ $\underset{\left(638\right)}{11.08}$ $\underset{\left(1203\right)}{11.17}$ $\underset{\left(2300\right)}{11.18}$ $\underset{\left(3657\right)}{11.18}$$\pm 0.05$
@@
Table 4: Outperformer spread option. The column labeled “GL” presents the results from [Guyon and Henry-Labordère, 2011] for the same option, while the column labeled “BM” shows the benchmark computed by using the change of numéraire technique. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
$N=32$ $N=64$ $N=128$
$\rho$ $P$ $125$ $250$ $500$ $1000$ $2000$ $125$ $250$ $500$ $1000$ $2000$ $125$ $250$ $500$ $1000$ $2000$ BM
$-0.5$ $\underset{\left(54\right)}{13.15}$ $\underset{\left(112\right)}{13.19}$ $\underset{\left(233\right)}{13.21}$ $\underset{\left(684\right)}{13.24}$ $\underset{\left(1892\right)}{13.25}$ $\underset{\left(101\right)}{13.27}$ $\underset{\left(210\right)}{13.29}$ $\underset{\left(456\right)}{13.31}$ $\underset{\left(1367\right)}{13.36}$ $\underset{\left(3725\right)}{13.36}$ $\underset{\left(188\right)}{13.31}$ $\underset{\left(393\right)}{13.28}$ $\underset{\left(935\right)}{13.32}$ $\underset{\left(2623\right)}{13.39}$ $\underset{\left(7090\right)}{13.40}$ $13.54$
00 $\underset{\left(54\right)}{13.09}$ $\underset{\left(100\right)}{13.12}$ $\underset{\left(226\right)}{13.18}$ $\underset{\left(739\right)}{13.22}$ $\underset{\left(1587\right)}{13.24}$ $\underset{\left(102\right)}{13.16}$ $\underset{\left(198\right)}{13.17}$ $\underset{\left(435\right)}{13.22}$ $\underset{\left(1291\right)}{13.30}$ $\underset{\left(3479\right)}{13.32}$ $\underset{\left(188\right)}{13.12}$ $\underset{\left(434\right)}{13.09}$ $\underset{\left(917\right)}{13.18}$ $\underset{\left(2701\right)}{13.28}$ $\underset{\left(7185\right)}{13.33}$ $13.52$
$0.5$ $\underset{\left(50\right)}{12.96}$ $\underset{\left(91\right)}{13.02}$ $\underset{\left(224\right)}{13.09}$ $\underset{\left(669\right)}{13.09}$ $\underset{\left(1312\right)}{13.11}$ $\underset{\left(96\right)}{12.95}$ $\underset{\left(181\right)}{13.00}$ $\underset{\left(410\right)}{13.08}$ $\underset{\left(1301\right)}{13.10}$ $\underset{\left(2874\right)}{13.14}$ $\underset{\left(191\right)}{12.90}$ $\underset{\left(364\right)}{12.91}$ $\underset{\left(878\right)}{13.01}$ $\underset{\left(2260\right)}{13.05}$ $\underset{\left(6707\right)}{13.11}$
NNU $N=32$ $N=64$ $N=128$
$\rho$ $E$ $100$ $200$ $400$ $800$ $1600$ $100$ $200$ $400$ $800$ $1600$ $100$ $200$ $400$ $800$ $1600$ ME BM
$-0.5$ $\underset{\left(72\right)}{12.21}$ $\underset{\left(130\right)}{12.93}$ $\underset{\left(245\right)}{13.03}$ $\underset{\left(476\right)}{13.05}$ $\underset{\left(939\right)}{13.05}$ $\underset{\left(129\right)}{12.30}$ $\underset{\left(230\right)}{13.17}$ $\underset{\left(433\right)}{13.30}$ $\underset{\left(839\right)}{13.32}$ $\underset{\left(1619\right)}{13.32}$ $\underset{\left(270\right)}{11.62}$ $\underset{\left(487\right)}{13.12}$ $\underset{\left(920\right)}{13.34}$ $\underset{\left(1786\right)}{13.38}$ $\underset{\left(3517\right)}{13.39}$$\pm 0.05$ $13.54$
00 $\underset{\left(73\right)}{12.30}$ $\underset{\left(130\right)}{13.03}$ $\underset{\left(247\right)}{13.15}$ $\underset{\left(479\right)}{13.18}$ $\underset{\left(944\right)}{13.19}$ $\underset{\left(147\right)}{12.49}$ $\underset{\left(266\right)}{13.21}$ $\underset{\left(504\right)}{13.33}$ $\underset{\left(980\right)}{13.36}$ $\underset{\left(1931\right)}{13.37}$ $\underset{\left(245\right)}{12.35}$ $\underset{\left(438\right)}{13.35}$ $\underset{\left(824\right)}{13.37}$ $\underset{\left(1596\right)}{13.41}$ $\underset{\left(3143\right)}{13.42}$$\pm 0.05$ $13.52$
$0.5$ $\underset{\left(84\right)}{11.64}$ $\underset{\left(153\right)}{12.79}$ $\underset{\left(294\right)}{12.96}$ $\underset{\left(574\right)}{13.00}$ $\underset{\left(1111\right)}{13.01}$ $\underset{\left(134\right)}{12.03}$ $\underset{\left(242\right)}{12.69}$ $\underset{\left(457\right)}{12.92}$ $\underset{\left(888\right)}{12.95}$ $\underset{\left(1749\right)}{12.95}$ $\underset{\left(248\right)}{11.99}$ $\underset{\left(442\right)}{12.97}$ $\underset{\left(831\right)}{13.17}$ $\underset{\left(1606\right)}{13.21}$ $\underset{\left(3159\right)}{13.21}$$\pm 0.05$
Table 5: Outperformer spread option with volatility bounds $\left[\sigma_{i}^{\min},\sigma_{i}^{\max}\right]=\left[0.1,0.4\right]$. The column labeled “BM” shows the benchmark computed by using the change of numéraire technique. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
@@ 5.1.3 Geo-Call spread
This case is a generalization of the UVM model in that the dividend is not constant but changes in relation to the value of volatility. The binomial tree method proposed by [Avellaneda et al., 1995] for the 1d UVM model can be used to calculate the price in this model, taking care to change the dividend in relation to volatility. In this regard, it should be noted that while in the one-dimensional UVM model the optimal $\sigma$ value is always $\sigma^{\min}$ or $\sigma^{\max}$ (this feature is known as “the bang bang condition”), in this model, characterized by a variable dividend yield $\hat{\eta}$, this property no longer applies. Therefore, at each node of the tree, the choice of optimal correlation is to be pursued not merely by comparing the objective function at the extreme values of $\sigma$, but by optimising the option price via an appropriate algorithm, such as the Sequential Quadratic Programming algorithm.
The results are shown in Table 6. As far as GTU results are concerned, the values obtained for varying $N$ and $P$ are very close to the benchmark for the considered values of $d$. The largest deviation is observed for $d=5$, with a difference of $0.04$ between the GTU price and the benchmark. However, this difference is minor, being less than $0.5\%$ of the benchmark. Additionally, we observe that the computational time increases linearly with the number of time steps $N$ and approximately linearly with the number of points $P$. This is particularly interesting since the cost for the GPR regression is cubic in $P$, indicating that most of the computational time is spent on solving optimization problems rather than on the regression itself. Lastly, the computational time grows exponentially with problem size, as the number of values to be evaluated with a tree step is $2^{d}$. Despite this, we are able to handle high dimensions, such as $d=10$, with reasonable computational effort.
The NNU results reported in Table 6 show excellent accuracy across all dimensions and discretisation parameters. For each value of $d\in\{2,5,10\}$, the estimated option price converges rapidly to the benchmark as the number of training epochs increases. The approximation is already close to the benchmark with just 200-400 epochs, and becomes essentially indistinguishable from it beyond 800 epochs. The results are particularly stable in higher dimensions, while computational times remain moderate, even for $d=10$, confirming the method’s scalability and robustness in multi-asset settings.
When compared with GTU, we observe that both methods achieve comparable levels of accuracy. However, NNU displays greater flexibility with respect to the dimension $d$, as its computational cost grows more moderately. In contrast, GTU becomes increasingly expensive for large $d$, due to the combinatorial growth of the tree. This highlights a key advantage of NNU in high-dimensional applications, where learning-based approaches offer a scalable alternative to traditional tree-based methods.
GTU $N$ $16$ $32$ $64$ BM
@@
NNU $N=16$ $N=32$ $N=64$
$d$ $E$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ $100$ $200$ $400$ $800$ ME BM
$2$ $\underset{\left(38\right)}{10.20}$ $\underset{\left(66\right)}{10.37}$ $\underset{\left(121\right)}{10.39}$ $\underset{\left(233\right)}{10.40}$ $\underset{\left(91\right)}{10.19}$ $\underset{\left(165\right)}{10.46}$ $\underset{\left(322\right)}{10.50}$ $\underset{\left(612\right)}{10.51}$ $\underset{\left(156\right)}{10.14}$ $\underset{\left(282\right)}{10.42}$ $\underset{\left(535\right)}{10.45}$ $\underset{\left(998\right)}{10.46}$$\pm 0.05$ $10.50$
$5$ $\underset{\left(47\right)}{9.58}$ $\underset{\left(84\right)}{9.65}$ $\underset{\left(161\right)}{9.67}$ $\underset{\left(314\right)}{9.68}$ $\underset{\left(82\right)}{9.56}$ $\underset{\left(147\right)}{9.62}$ $\underset{\left(278\right)}{9.66}$ $\underset{\left(539\right)}{9.68}$ $\underset{\left(211\right)}{9.54}$ $\underset{\left(387\right)}{9.63}$ $\underset{\left(734\right)}{9.65}$ $\underset{\left(1369\right)}{9.66}$$\pm 0.03$ $9.70$
$10$ $\underset{\left(47\right)}{9.49}$ $\underset{\left(82\right)}{9.54}$ $\underset{\left(152\right)}{9.55}$ $\underset{\left(309\right)}{9.55}$ $\underset{\left(92\right)}{9.46}$ $\underset{\left(171\right)}{9.54}$ $\underset{\left(329\right)}{9.54}$ $\underset{\left(645\right)}{9.54}$ $\underset{\left(195\right)}{9.45}$ $\underset{\left(358\right)}{9.52}$ $\underset{\left(678\right)}{9.53}$ $\underset{\left(1310\right)}{9.53}$$\pm 0.02$ $9.55$
Table 6: Geo-Call spread, with $\rho_{i,j}=0$ and with $M=2^{d}$. The column labeled “BM” shows the benchmark obtained via one-dimensional reduction. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
Table 7 reports the performance of GTU and NNU in the pricing of a high-dimensional geometric call spread under zero correlation. To mitigate the exponential growth in GTU’s computational cost as the dimension $d$ increases, we adopt a sampling-based strategy: instead of summing over all $2^{d}$ possible branches, we randomly select a subset of $M$ representative terms. This Monte Carlo-type simplification is theoretically justified, as the approximated price converges to the full-sum value as $M\to 2^{d}$. Empirically, the results in Table 7 show that even a few thousand samples are sufficient to obtain accurate and stable estimates, enabling the use of GTU for dimensions as high as $d=40$ within reasonable computation times. Nevertheless, the GTU approach becomes significantly more costly and less effective for very large dimensions, such as $d=80$. In this regime, we observe both a substantial increase in computational burden and a slower convergence to the benchmark, particularly for coarse discretizations. This illustrates a key limitation of tree-based methods, which are inherently affected by the curse of dimensionality.
By contrast, the NNU method exhibits outstanding stability and accuracy across all tested dimensions. It consistently produces prices that are virtually indistinguishable from the benchmark, with negligible margins of error and computational times that grow only linearly with the number of training epochs. Even for $d=80$, NNU delivers precise results with manageable computational effort, highlighting its scalability and robustness.
In conclusion, while GTU remains effective in low- to moderate-dimensional settings, its performance deteriorates as the dimensionality grows. NNU clearly outperforms GTU in high dimensions, offering a more efficient and practical solution for the valuation of derivatives in large-scale market models.
GTU $N=16$ $N=32$ $N=64$